# CMRNext: Camera to LiDAR Matching in the Wild for Localization and Extrinsic Calibration

### v4

@@ IV-K Extrinsic Calibration — Generalization
IV-L Hyperparameter Optimization
The default hyperparameter values used for all the experiments reported in previous sections were optimized using the hyperparameter optimization method BOHB [71], which combines Bayesian optimization and Hyperband. In particular, we used the implementation provided with the Weights&Biases platform [72]. The hyperparameters optimized were the learning rate, the batch size, the number of Fourier frequencies used for the positional encoding, the parameters of the occlusion filter, the optimizer, the learning rate scheduler, and the maximum depth at which a point in the LiDAR image is kept. Table X shows the search space used for each hyperparameter and the optimal value found by the hyperparameter optimization.
@@ IV-L Hyperparameter Optimization
IV-M Runtime Analysis
In this experiment, we compare the runtime of CMRNext with existing camera-agnostic methods CMRNet++ [27] and I2D-Loc [65]. Additionally, we compare the performance and runtime of different algorithms to solve the PnP problem, most of which are implemented in the OpenCV library [73], including EPnP [50], SQPnP [74], and our GPU-based implementation of EPnP + RANSAC. All runtimes were computed only for the first iteration on a workstation with 32GB DDR3 RAM, an Intel i7-4790K CPU, and an NVIDIA RTX 2080 Ti GPU. For the iterative refinement approach described in Section III-E, the runtime scales linearly with the number of iterations. The results reported in Table XI show that CMRNext with the GPU-based PnP solver achieves the fastest runtime among all methods. Moreover, it is possible to further improve the performance of CMRNext by performing a Levenberg–Marquardt (LM) optimization on the inliers found by RANSAC at the cost of increased runtime.
@@ IV-M Runtime Analysis
V Limitations
CMRNext has three primary limitations. First, because it is trained exclusively on autonomous driving datasets, its generalization ability is likely to be limited in other environments, such as off-road or mining environments. Second, while our method is faster than existing camera-agnostic approaches, it is not yet capable of real-time performance. Lastly, similar to most existing methods, the training paradigm requires ground truth poses as well as correct intrinsic camera parameters, which are typically provided in publicly available datasets, but add an additional constraint to the approach.
VI Conclusions
In this paper, we presented CMRNext for monocular localization in LiDAR-maps and extrinsic LiDAR-camera calibration in the wild. CMRNext decomposes the LiDAR-camera relative pose estimation problem in two steps. First, an optical flow estimation network predicts dense point-pixel correspondences between the camera image and the point cloud. Subsequently, we use these matches to predict the relative pose between the camera and the LiDAR using robust geometric techniques. Through extensive experimental evaluations, we demonstrated that CMRNext sets the new state-of-the-art on both monocular localization in LiDAR-maps and extrinsic calibration tasks. Remarkably, CMRNext can effectively generalize without any retraining or fine-tuning to different environments and different robotic platforms, including three in-house robots with diverse locomotion: a self-driving perception car, a quadruped robot, and a quadcopter UAV. CMRNext achieves median translation and rotation errors of and , respectively, for the localization task on the sequence 00 of the KITTI odometry dataset. Additionally, CMRNext calibrated the LiDAR-camera extrinsic parameters of our in-house platforms with translation and rotation errors as low as and , respectively, demonstrating exceptional zero-shot generalization ability. To foster research in this direction, we have made the code and pre-trained models publicly available at http://cmrnext.cs.uni-freiburg.de.

### v5

@@ IV-L Extrinsic Calibration — Generalization
IV-M Hyperparameter Optimization
The default hyperparameter values used for all the experiments reported in previous sections were optimized using the hyperparameter optimization method BOHB [80], which combines Bayesian optimization and Hyperband. In particular, we used the implementation provided with the Weights&Biases platform [81]. The hyperparameters optimized were the learning rate, the batch size, the number of Fourier frequencies used for the positional encoding, the parameters of the occlusion filter, the optimizer, the learning rate scheduler, and the maximum depth at which a point in the LiDAR image is kept. Table XI shows the search space used for each hyperparameter and the optimal value found by the hyperparameter optimization.
@@ IV-M Hyperparameter Optimization
IV-N Runtime Analysis
In this experiment, we compare the runtime of CMRNext with existing camera-agnostic methods CMRNet++ [30] and I2D-Loc [65]. Additionally, we compare the performance and runtime of different algorithms to solve the PnP problem, most of which are implemented in the OpenCV library [82], including EPnP [49], SQPnP [83], and our GPU-based implementation of EPnP + RANSAC. All runtimes were computed only for the first iteration on a workstation with 32GB DDR3 RAM, an Intel i7-4790K CPU, and an NVIDIA RTX 2080 Ti GPU. For the iterative refinement approach described in Section III-E, the runtime scales linearly with the number of iterations. The results reported in Table XII show that CMRNext with the GPU-based PnP solver achieves the fastest runtime among all methods. Moreover, it is possible to further improve the performance of CMRNext by performing a Levenberg–Marquardt (LM) optimization on the inliers found by RANSAC at the cost of increased runtime.
@@ IV-N Runtime Analysis
V Conclusions and Limitations
In this paper, we presented CMRNext for monocular localization in LiDAR-maps and extrinsic LiDAR-camera calibration in the wild. CMRNext decomposes the LiDAR-camera relative pose estimation problem in two steps. First, an optical flow estimation network predicts dense point-pixel correspondences between the camera image and the point cloud. Subsequently, we use these matches to predict the relative pose between the camera and the LiDAR using robust geometric techniques. CMRNext has three primary limitations. First, since it is trained exclusively on autonomous driving datasets, its generalization ability is likely to be limited in other environments, such as off-road or mining environments. Second, while our method is faster than existing camera-agnostic approaches, it is not yet capable of real-time performance. Lastly, similar to most existing methods, the training paradigm requires ground truth poses as well as correct intrinsic camera parameters, which are typically provided in publicly available datasets, but add an additional constraint to the approach. Through extensive experimental evaluations, we demonstrated that CMRNext sets the new state-of-the-art on both monocular localization in LiDAR-maps and extrinsic calibration tasks. Remarkably, CMRNext can effectively generalize without any retraining or fine-tuning to different environments and different robotic platforms, including three in-house robots with diverse locomotion: a self-driving perception car, a quadruped robot, and a quadcopter UAV. CMRNext achieves median translation and rotation errors of and , respectively, for the localization task on the sequence 00 of the KITTI odometry dataset. Additionally, CMRNext calibrated the LiDAR-camera extrinsic parameters of our in-house platforms with translation and rotation errors as low as and , respectively, demonstrating exceptional zero-shot generalization ability. To foster research in this direction, we have made the code and pre-trained models publicly available at http://cmrnext.cs.uni-freiburg.de.